# UniCode: A Framework for Generating High-Quality Competitive Coding Problems

### v1

@@ Abstract
The reliance of competitive coding benchmarks on static, human-authored problems creates significant challenges, including data contamination and limited scalability. To address these issues, we introduce UniCode, a novel framework that automatically generates high-quality algorithmic problems alongside robust, contamination-resistant test cases. Inspired by biological evolution that creates better and diverse offspring, our framework leverages Large Language Models (LLMs) to systematically diversify problems through three strategies: single-problem extension, same-type fusion, and cross-type fusion. A key innovation is our stress-driven test case synthesis pipeline, which generates reliable test suites without requiring a canonical ground-truth solution. This pipeline combines brute-force grounding for small-scale inputs with a consensus-based validation mechanism for large-scale inputs to ensure high correctness and coverage. We demonstrate effectiveness of our framework by curating a benchmark of 492 problems and evaluating 19 state-of-the-art LLMs. The results reveal that UniCode is highly challenging and discriminative, with the top-performing model, o4-mini, achieving a pass rate of only 70.3%. Our framework provides a scalable and reliable solution for generating dynamic evaluation datasets in coding domain.
@@ 1 Introduction
The rapid advancement of large language models (LLMs) in code generation (Li et al., 2022; Allal et al., 2023; Jaech et al., 2024; Guo et al., 2024; Hui et al., 2024; Comanici et al., 2025; Guo et al., 2025; Zhao et al., 2025) presents an evaluation paradox. While models are approaching saturation on existing static benchmarks (Zhu et al., 2025; Chen et al., 2021; Austin et al., 2021; Hendrycks et al., 2021), it is increasingly uncertain whether this performance reflects genuine reasoning ability or simply vast memorization of training data (Oren et al., 2023; Roberts et al., 2023; Golchin & Surdeanu, 2023; Riddell et al., 2024; Tang et al., 2024; Li et al., 2025). Even advanced benchmarks that periodically add new problems, like LiveCodeBench (Jain et al., 2024) and LiveCodeBench-Pro (Zheng et al., 2025b), are constrained by their reliance on costly, slow human authoring, which limits their scalability. This core issue is rendering static evaluation obsolete and demands a fundamental shift in methodology. We argue that the next frontier lies in generative evaluation (Zhu et al., 2023; Lin et al., 2025; Parmar et al., 2024; Shi et al., 2025; Zheng et al., 2025a): a paradigm where the benchmark itself is an intelligent system, dynamically producing a near-endless stream of novel challenges to probe the true generalization capabilities of models.
To meet this challenge, we introduce UniCode, a generative evaluation framework that automatically synthesizes novel, competition-level programming problems and robust test suites at scale. At the heart of our framework is a problem generation engine inspired by biological evolution (Kutschera & Niklas, 2004; Pagel, 1999; Koonin, 2011). The diversity of life is driven by two key mechanisms: gene mutation and genetic recombination (Baake & Gabriel, 2000; Charlesworth et al., 2009). Similarly, UniCode employs three complementary strategies to create new tasks from a seed set (as shown in Figure˜5): single-problem extension, which modifies a problem to increase its complexity (e.g., evolving Two Sum to Three Sum); same-type fusion, which combines two problems with the same algorithmic tag to create a new variant with similar logic but a different narrative; and cross-type fusion, which merges problems from different categories to form complex hybrid challenges (e.g., Palindrome Sum Pair). This methodology can produce well-formed and structurally novel problems, with a solvability rate of 98.0% confirmed by blind human review.
A key innovation of UniCode is its stress-driven test case synthesis pipeline (Figure˜2), which ensures the quality and reliability of the evaluation without needing a canonical ground-truth solution. This process first establishes trusted outputs for small-scale inputs using a brute-force solver, which are then used to “stress-test” and filter a pool of LLM-generated optimized solutions. For large-scale inputs, where brute force is infeasible, correctness is determined by a majority-vote consensus among the filtered, high-quality solvers, with a powerful LLM adjudicating disputes to enhance reliability. Our experiments demonstrate that this pipeline achieves a test case correctness rate of 94.5% while also providing robust coverage that effectively identifies flaws in imperfect solutions.
We demonstrate our framework’s effectiveness by curating a new benchmark of 492 problems and evaluating 19 state-of-the-art LLMs. Our results confirm that UniCode is highly challenging and discriminative: the top-performing model, o4-mini, achieves a pass@1 rate of only 66.1%. A controlled experiment validates that our generation strategy effectively creates new problems that are more challenging, as models performance drops on average by over 30% on UniCode problems compared to negligible drops on simple narrative paraphrases for seed problems(see Figure˜3). The evaluation also reveals that leading open-source models are highly competitive, signaling a promising direction for reproducible and accessible AI research. Insummary, UniCode provides a scalable, reliable, and contamination-resistant solution for the dynamic evaluation of competitive coding.
2 Methodology
This section details the UniCode framework, which systematically generates novel competitive programming problems and their corresponding robust test suites. Our approach is divided into two core pipelines: problem generation (Section 2.1) and test case generation (Section 2.2).
2.1 Problem generation
Figure 1: An illustration of our three problem generation strategies using seed problems. Same-type fusion (left): combines two problems sharing a common tag (e.g., hash table) to produce a new problem with the same core logic but greatly different descriptions. Single-problem extension (middle): modifies a seed problem to create a more complex version (e.g., Two Sum $\rightarrow$ Three sum). Cross-type fusion (right): merges problems from different categories to create a hybrid challenge (e.g., Palindrome Sum Pair). For concrete examples of generated problems, please refer to Appendix appendix˜F.
The key idea behind generating new problems is to modify existing ones. We first observe that naive superficial changes to problems, such as changing the background story or adding distracting information, do not produce challenging or novel tasks, as confirmed by our experiments (Figure˜3). Therefore, we design generative process that are more powerful to create diverse problems for our framework. Let $\mathcal{D}$ be a dataset of programming problems, where each problem is tagged with one or more algorithm types from a set $\mathcal{T}$. To generate a new problem, $p_{\mathrm{new}}$, we sample one or two seed problems from $\mathcal{D}$ and apply one of three strategies:
Single-problem extension
This strategy creates a new problem by modifying a single seed problem $p_{a}$. We provide an LLM with $p_{a}$ and a modification instruction $\phi$. These instructions can include actions like tightening constraints, altering the input format, or adding a side condition. The LLM then generates a new problem $p_{\mathrm{new}}$, that retains the core structure of the original but introduces a fresh challenge. For example, the LLM can evolve the classic Two Sum problem into Three sum, a more complex variant requiring a different algorithmic approach (Figure˜5, bottom middle).
Same-type fusion
This strategy generates a novel problem by fusing two seed problems $p_{a}$ and $p_{b}$, that share a common algorithmic tag $t$ (e.g., hash-table). The LLM is prompted to first identify the underlying solution pattern shared by both problems and then instantiate this pattern into a new problem $p_{\mathrm{new}}$, with a distinct narrative but the same core solution logic. For example, by analyzing two different problems that both use hash tables for efficient lookups, the model can generate a new frequency-counting task that leverages the same core data structure (Figure˜5, bottom left). Providing concrete seed problems as in-context examples, rather than just a tag $t$, is crucial. It helps constrain the LLM to generate new problems of comparable quality and difficulty, which leads to a greater diversity of more appropriate challenges.
Cross-type fusion
This strategy creates more difficult, hybrid problems that require integrating multiple skills. We start with two seed problems $p_{a}$ and $p_{b}$, from different algorithmic categories. The LLM identifies a “bridging concept”—a compatible element that can link the two problems. This concept is then used to integrate a feature from one problem into the other, generating a composite problem $p_{\mathrm{new}}$. For example, merging a Two Sum problem (hash table) with a Palindrome checking problem (strings, dynamic programming) could result in a new Palindrome Sum Pair task, which requires a solution that combines both skills (Figure˜5 bottom right).
2.2 Test Case Generation
Generating reliable test cases for novel problems without a known ground-truth solution is a significant challenge. We address this with a two-part pipeline that first generates diverse inputs and then establishes trusted outputs through a hybrid brute-force and consensus-based approach.
2.2.1 Input Generation
@@
Based on previous work in test case generation (Jain et al., 2024; Liu et al., 2025; Wang et al., 2025), we construct a comprehensive test suite for each problem by prompting LLMs to generate inputs from three complementary sources:
  1. 1.
@@
    Random Generation ($G_{\mathrm{rand}}$): A Python function that samples broadly from the valid input space to ensure wide coverage, including both small-cale and large-scale instances.
  2. 2.
    Adversarial Generation ($G_{\mathrm{adv}}$): A Python function that targets boundary conditions and worst-case scenarios using predefined strategies (e.g., maximum values, alternating values) to probe algorithmic robustness and efficiency.
  3. 3.
    LLM-based Synthesis ($G_{\mathrm{llm}}$): This method synthesizes full input text, offering greater flexibility without external libraries. It produces small-scale yet challenging inputs designed to expose subtle, hard-to-find failure modes, and is categorized as adversarial in our experiments.
@@
All candidate inputs, regardless of their source, are validated by the verifier $V$. For each input $\tau\in\{\mathrm{rand},\mathrm{adv},\mathrm{llm}\}$, let $G_{\tau}$ be the candidate pool. The final set of verified inputs $I_{\tau}$ is defined as:
@@
To balance coverage and efficiency (Liu et al., 2023a), we assemble a final test suite $S$ of 50 cases with a fixed composition: 20 random, 20 adversarial, and 10 LLM-synthesized inputs. We discuss in Table˜1 to show this setting has good correctness and coverage.
2.2.2 Output Generation
Figure 2: Stress-driven pipeline for ground-truth generation. (1) Stress Testing: A brute-force solver establishes trusted outputs for small-scale inputs. These are used as a "stress test" to filter a pool of LLM-generated optimized solutions. (2) Consensus Validation: The filtered, efficient solvers are run on large-scale inputs, and outputs are determined by a strict majority vote. (3) LLM Adjudication: A powerful LLM adjudicates between conflicting outputs for inputs where no majority is reached.

### v2

# UniCode: Augmenting Evaluation for Code Reasoning

@@ Abstract
Current coding benchmarks often inflate Large Language Model (LLM) capabilities due to static paradigms and data contamination, enabling models to exploit statistical shortcuts rather than genuine reasoning. To address this, we introduce UniCode, a generative evaluation framework that systematically probes LLM limits via: (1) multi-dimensional augmentation transforming seed problems into complex variations to disrupt fixed algorithmic patterns; (2) a highly reliable, automated test generation pipeline for scalable evaluation; and (3) fine-grained metrics for rich error signals. Experiments reveal a 31.2% performance collapse in state-of-the-art models on UniCode, primarily driven by deficiencies in conceptual modeling and scalability reasoning rather than syntactic errors. Furthermore, we uncover a “seed-problem regression” where models revert to memorized seed logic rather than following new specifications, signaling a reliance on shortcuts over reasoning. This work validates UniCode as a robust framework to expose model fragility and foster reasoning-oriented code intelligence.
Machine Learning, ICML
@@ 1 Introduction
Developing intelligent systems capable of multi-step reasoning remains a cornerstone of AI research (Wei et al., 2022; Guo et al., 2025; Jaech et al., 2024; Comanici et al., 2025; Li et al., 2022a). Competitive programming has emerged as an ideal testbed for evaluating such capabilities (Li et al., 2022c; El-Kishky et al., 2025), not merely for its rigorous evaluation signals, but because it establishes coding as a general-purpose interface for problem-solving. In this context, writing executable code becomes a universal medium for grounding complex reasoning and formalizing solutions across diverse tasks (Zhu et al., 2025; Quan et al., 2025).
However, current benchmarks suffer from an “evaluation paradox”: while LLMs have nearly achieved saturation on standard coding benchmarks (Chen et al., 2021; Austin et al., 2021; Hendrycks et al., 2021), they frequently stumble during real-world interactions (Sergeyuk et al., 2025; Weisz et al., 2025). We attribute this discrepancy to three critical limitations in existing evaluation protocols: 1) data contamination and fixed design patterns, which allow models to exploit statistical shortcuts (Figure 1 (d)); 2) limited scalability due to the high cost of human curation (Jain et al., 2024); and 3) a reliance on static datasets that fail to capture the open-ended nature of real-world software engineering (Fodor, 2025; Stanley et al., 2019).
To address these issues, recent research has explored dataset augmentation via perturbation (Li et al., 2024a; Mirzadeh et al., 2024; Orvalho and Kwiatkowska, 2025). Yet, these approaches predominantly focus on superficial variations, such as variable renaming or surface-level rephrasing, rather than altering the underlying reasoning logic. Consequently, they fail to assess whether a model has truly mastered the algorithmic concepts or is merely recalling specific problem structures. To rigorously evaluate adaptability, there is an urgent need for a systematic augmentation framework that mimics real-world challenges through deep, structural transformations rather than shallow manipulations.
In this work, we introduce UniCode, a framework for the Augmented Evaluation of code reasoning (Figure 1 (a)), which employs a generative approach to systematically stress-test models under meaningful structural, compositional, and conceptual shifts. Our framework offers the following three key contributions:
Systematic Task Augmentation. We propose a rigorous methodology that transforms high-quality “seed problems” into a diverse array of challenging tasks in Section 2. Specifically, Atomic variations modify targeted task facets (e.g., narrative, rules, or input scale) to determine if models can adapt to structural changes rather than relying on rote retrieval; Compositional variations integrate multiple knowledge points to generate novel combinatorial patterns, forcing models to engage in genuine multi-step reasoning (see  Figure 1 (b)). By applying broadly meaningful transformations ranging from narrative rewriting to logical reconfiguration, we simulate the complexity of real-world requirements.
Scalable and Robust Evaluation. To address the scalability bottleneck of human curation, we develop a stress-driven synthesis pipeline in Section 3. By integrating stress-testing with multi-model consensus validation, our approach achieves 94.5% reliability in test case generation without requiring ground-truth solutions. Test cases from three distinct sources: random sampling, adversarial boundary cases, and LLM-synthesized hard examples, guarantee a robust assessment of functional correctness and corner-case stability. This allows for the dynamic expansion of an effectively infinite, contamination-resistant problem space, ensuring the benchmark remains challenging as models evolve.
Fine-grained Diagnostic Metrics. Current evaluations often rely on a single 0-1 pass rate, which obscures the specific deficiencies of the model. UniCode transcends this limitation by providing a comprehensive diagnostic toolkit. We categorize errors into modeling flaws, logic gaps, algorithmic complexity misjudgments, and implementation bugs in Section 5.2. This granular analysis moves beyond simple accuracy to reveal why a model fails, offering practical insights for improving code reasoning capabilities.
We benchmark 19 LLMs on UniCode, finding it highly challenging and discriminative. Even the top-performing model, o4-mini, achieves a pass rate of only 70.3%. Our in-depth analysis reveals a critical vulnerability: while LLMs remain robust against superficial perturbations, their performance collapses ($30\%$) under our structural augmentations. Case studies reveal that models often remain “trapped” in the reasoning paradigms of the seed problems, failing to adapt to the augmented logic (Figure 5). By providing a fine-grained view of these failures, UniCode paves the way for the development of more robust, generalized multi-step code agents.
Figure 1: An Overview of UniCode. (a) UniCode Framework. We propose a generative pipeline to create a vast evaluation space via multi-dimensional augmentations to probe reasoning axes, a test generation module for scalability, and fine-grained diagnostics that expose specific failure modes for a transparent assessment of code reasoning. (b) An Example: Probing Code Reasoning Ability via Five Augmentation Axes. We design Atomic Augmentations (narrative, rule, and efficiency) to alter task facets, ranging from surface-level shifts to reconfigurations of the reasoning graph topology; and Compositional Augmentations (sequential composition and concept fusion) to disrupt fixed algorithmic patterns and test combinatorial generalization. (c) Model Fragility: Vulnerability to Logic Alteration. While LLMs are robust under surface-level narrative shifts, performance drops significantly when core logic is modified. The most severe failures in sequential and fusion tasks indicate struggles with long causal chains and algorithmic integration. (d) Beyond Human Curation: Identifying Reasoning Blind Spots. Unlike static benchmarks (e.g., TACO (Li et al., )) with predictable patterns, UniCode explores a broader spectrum of algorithmic combinations to expose “reasoning blind spots” that human designers typically fail to cover. (e) Performance Collapse: Memorization vs. Adaptation. Large-scale testing reveals an average drop of $30\%$ when moving from seed problems to their augmented variations across LLMs. We observe a “seed-problem regression” where models default to the memorized logic of seed problems rather than adapting to new, augmented constraints. Best viewed when zoom in.
Figure 2: Stress-driven pipeline for ground-truth generation. (1) Stress Testing: we use a brute-force solver to produce trusted outputs on small inputs, which serve as stress tests to filter LLM-generated optimized solvers. (2) Consensus Validation: The remaining solvers are executed on large inputs, and the output is selected by strict majority vote. (3) LLM Adjudication: A powerful LLM adjudicates between conflicting outputs for inputs where no majority is reached. The ablation study in App. B.1 verifies the effectiveness of each stage.
2 Augmentation Axes for Code Reasoning
Data-driven LLMs often rely on statistical correlations over logical reasoning, faltering in complex, open-world scenarios (McCoy et al., 2019). While perturbations are standard for testing, existing benchmarks struggle to challenge increasingly robust models.To address this, we design Atomic Augmentations (narrative, rule, and efficiency) ranging from surface shifts to structural reasoning graph reconfigurations; and Compositional Augmentations (sequential composition and concept fusion) to disrupt fixed algorithmic patterns and evaluate combinatorial generalization. These methods are illustrated in Figure 1 (b) and detailed below:
Narrative Perturbation
This axis modifies variable names, thematic backgrounds, or injects irrelevant contextual noise without altering the underlying logic. It specifically probes whether a model suffers from “token bias” (Jiang et al., 2024) or exhibits content-agnostic reasoning. For example, we reframe the abstract Longest Increasing Subsequence (LIS) problem into a real-world scenario like Identifying the Longest Growth Period in Stock Trends (see Figure 1 (b)). If a model fails due to this narrative shift, it indicates a reliance on near-neighbor matching (Li et al., 2024b) of familiar problem descriptions rather than a robust understanding of the logical core.
Rule Modification
Standard programming problems often have “canonical” solutions that LLMs easily memorize. By subtly altering operational rules or boundary conditions, we invalidate these memorized paths (Dziri et al., 2023). A representative transformation is shifting LIS to the Longest Non-Decreasing Subsequence. While seemingly minor, this change shifts the comparison operator (from >> to $\geq$), requiring the model to re-calibrate its logical flow according to novel instructions rather than retrieving pre-trained code snippets, effectively distinguishing retrieval from reasoning.
Efficiency Scaling
Genuine reasoning entails an awareness of the computational budget and the ability to adapt as data scale increases (Zubić et al., 2025). This axis tests if a model can transition from a naive approach to a more optimized algorithm when complexity demands it. For instance, when the input size $n$ for LIS scales to $10^{5}$, the standard $O(n^{2})$ approach becomes computationally prohibitive. The model must recognize this bottleneck and pivot to a greedy strategy ($O(n\log n)$). This transition probes the model’s capacity for high-level strategy selection and its understanding of algorithmic efficiency beyond simple template filling.
Sequential Composition
This dimension involves chaining multiple distinct algorithmic steps to examine the stability of the reasoning chain. As the sequence of operations grows, the probability of failure increases—a phenomenon known as error propagation (Schaeffer et al., 2023). In a composite variant like the Longest Bitonic Subsequence, the model must compute the LIS from both the prefix and suffix and then integrate the results. Such tasks reveal the fragility of the reasoning process, as minor logical flaws that might be hidden in single-step tasks are amplified during intermediate state transfers.
Concept Fusion
Real-world challenges often lie at the intersection of disparate domains. The fusion variant merges distinct algorithmic concepts into a single problem, creating novel combination patterns that human designers often overlook. For instance, while dynamic programming is frequently paired with string manipulation in standard datasets, it is rarely combined with game theory or greedy algorithms, making such intersections particularly difficult to navigate. By creating these original pairings, we probe whether the model can genuinely integrate separate concepts to achieve combinatorial generalization, which is recognized as a cornerstone of human-like reasoning (Battaglia et al., 2018).
By evaluating performance across these strategic axes that reconfigure the reasoning graph topology, we systematically probe the model’s structural generalization and its adaptability to novel coding challenges.
3 Scalable and Rigorous Test Generation
To scale evaluation effectively, an automated test generation pipeline is essential. However, the core challenge is ensuring test quality for novel problems lacking human-authored solutions. We address this by using three distinct input types to cover boundary conditions and attacks, then establishing trusted outputs through a stress-driven pipeline.
3.1 Input Generation
@@
Low-quality test cases often lead to model mis-ranking (Jain et al., 2024; Liu et al., 2025; Wang et al., 2025). To ensure robustness, we construct test cases by prompting LLMs to generate inputs from three complementary sources:
@@
  • Random Generation ($G_{\mathrm{rand}}$): targets general correctness that samples broadly from the valid input space.
  • Adversarial Generation ($G_{\mathrm{adv}}$): targets algorithmic inefficiencies or edge-case failures, such as boundary-value, extreme sequence lengths, and oscillating patterns.
  • Corner Generation ($G_{\mathrm{corn}}$): targets subtle failure modes using challenging and small-scale inputs.
@@
All candidate inputs from each source $\tau\in\{\text{rand, adv, corn}\}$ (input prompts detailed in App. Section B.7) are filtered by a verifier $V$. Let $G_{\tau}$ represent the initial candidate pool; the final set of verified inputs $I_{\tau}$ is defined as:
@@
To balance coverage and efficiency (Liu et al., 2023a), we specifically assemble a test suite $S$ of diverse 50 cases: 20 random, 20 adversarial, and 10 corner cases (empirically tuned; see App. Section B.2). Table 3 demonstrates that this configuration ensures both correctness and coverage.
Model Difficulty (Pass@1) Test Impact ($\Delta\downarrow$) Avg. Pass@1 Cost / Prob. (\$)
Easy Medium Hard $\Delta_{Rand}$ $\Delta_{Adv}$ $\Delta_{Corn}$
Reasoning Models
o4-mini (high)* 94.9% 78.2% 21.6% $\mathit{-2.5\%}$ $\mathit{-6.6\%}$ $\mathit{-10.2\%}$ 70.3% 0.0269
gpt-5 (medium)* 89.5% 77.6% 18.8% $\mathit{-2.3\%}$ $\mathit{-5.6\%}$ $\mathit{-8.9\%}$ 67.7% 0.0390
o4-mini (medium)* 89.2% 73.6% 20.3% $\mathit{-2.7\%}$ $\mathit{-8.8\%}$ $\mathit{-11.8\%}$ 66.1% 0.0205
google/gemini-2.5-pro* 94.0% 53.1% 8.5% $\mathit{-2.8\%}$ $\mathit{-7.2\%}$ $\mathit{-9.3\%}$ 61.6% 0.2015
deepseek-v3.1 (thinking) 89.2% 59.8% 11.5% $\mathit{-2.0\%}$ $\mathit{-4.1\%}$ $\mathit{-8.1\%}$ 60.5% 0.0276
deepseek-r1 80.3% 36.4% 5.1% $\mathit{-3.2\%}$ $\mathit{-6.8\%}$ $\mathit{-7.3\%}$ 55.6% 0.0250
o3-mini (medium)* 86.2% 50.0% 6.0% $\mathit{-2.8\%}$ $\mathit{-7.8\%}$ $\mathit{-7.0\%}$ 55.1% 0.0230
qwen3-235b-a22b 80.2% 39.7% 5.1% $\mathit{-2.3\%}$ $\mathit{-9.9\%}$ $\mathit{-13.3\%}$ 53.5% 0.0343
gemini-2.5-flash* 81.4% 22.6% 4.8% $\mathit{-2.6\%}$ $\mathit{-7.7\%}$ $\mathit{-6.3\%}$ 47.7% 0.0090
grok-3-mini* 77.8% 21.7% 3.3% $\mathit{-2.1\%}$ $\mathit{-10.0\%}$ $\mathit{-6.4\%}$ 46.4% 0.0035
claude-3.7-sonnet* 76.2% 24.1% 2.4% $\mathit{-1.2\%}$ $\mathit{-11.5\%}$ $\mathit{-5.2\%}$ 45.5% 0.1282
Non-Reasoning Models
deepseek-chat-v3.1 82.7% 29.3% 3.9% $\mathit{-2.0\%}$ $\mathit{-5.1\%}$ $\mathit{-7.3\%}$ 49.8% 0.0068
gpt-4.1-mini* 73.7% 20.9% 3.8% $\mathit{-2.8\%}$ $\mathit{-8.8\%}$ $\mathit{-7.4\%}$ 42.4% 0.0070
gpt-4.1* 62.1% 21.8% 1.4% $\mathit{-3.0\%}$ $\mathit{-10.2\%}$ $\mathit{-10.4\%}$ 36.5% 0.0071
qwen3-coder 66.5% 9.3% 0.0% $\mathit{-1.3\%}$ $\mathit{-10.5\%}$ $\mathit{-9.9\%}$ 35.4% 0.0145
claude-sonnet-4* 60.7% 14.0% 2.0% $\mathit{-1.5\%}$ $\mathit{-13.1\%}$ $\mathit{-5.1\%}$ 32.4% 0.0211
llama-4-maverick 51.3% 8.6% 0.0% $\mathit{-1.3\%}$ $\mathit{-9.5\%}$ $\mathit{-8.8\%}$ 26.2% 0.0006
gpt-4o* 31.3% 2.2% 0.0% $\mathit{-0.4\%}$ $\mathit{-2.2\%}$ $\mathit{-6.4\%}$ 15.4% 0.0139
qwen-2.5-32b-coder 27.2% 2.2% 0.0% $\mathit{-1.8\%}$ $\mathit{-0.3\%}$ $\mathit{-5.5\%}$ 13.4% 0.0038
gemma-3-27b-it 26.1% 2.2% 0.0% $\mathit{-0.2\%}$ $\mathit{-4.3\%}$ $\mathit{-3.5\%}$ 13.1% -
llama-3.3-8b-instruct 11.2% 1.1% 0.0% $\mathit{-0.4\%}$ $\mathit{-0.2\%}$ $\mathit{-0.3\%}$ 5.5% 0.0002
Table 1: UniCode Leaderboard. We report Pass@1 across three difficulty levels and evaluate model robustness via $\Delta$ ($\Delta=\text{Overall}-\text{w/o Test}$), where larger drops indicate greater vulnerability. The significant drops ($\Delta\downarrow$) demonstrate the UniCode’s capability to expose reasoning flaws. The cost per problem is included to help identify cost-effective models for future research. (*: closed models).
3.2 Ground-Truth Construction